### v13

@@ 6 Checking the mecanism optimality for each generation
We observe first that if (9) holds, then $V$ defined in (14) satisfies the Bellman equation in (15). Indeed, since by definition of $\tau^{*}$ we have $u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))=\overline{v}_{1}(\lambda_{t})$,
$V(\lambda_{t},t)$ $=\mathrm{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathrm{E}_{t}\big{[}\overline{v}_{1}(\lambda_{t+1})|\overline{v}_{1}(\lambda_{t})\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta V(\lambda_{t+1},t+1).$ (16)
The last equation follows from the fact that $\lambda_{t+1}$ is realized given $e_{t}$ and seeing $\mathrm{E}_{t+1}$ as a function of $e_{t+1}$. But $\tau^{*}$ is the solution to the problem
$\sup_{\tau\in\Lambda(\lambda_{t},t)}(1-\beta)u(e_{t}+\tau(e_{t}))+\beta\mathrm{E}_{t}\big{[}\overline{v}_{1}(\lambda_{t+1})|\overline{v}_{1}(\lambda_{t})\big{]}$ (17)
under the binding constraint
$(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k}\tau(e_{k})\right]=\mathrm{E}[\tau(e_{t})]=\text{Const.}$. (18)
@@
Indeed, for a given $\lambda_{t}$ the objective function in (17) multiplied by $\lambda_{t}$ is written as
$\lambda_{t}\big{[}(1-\beta)$ $u(e_{t}+\tau(e_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}$
$=\lambda_{t}\big{[}(1-\beta)u(e_{t}+\tau(e_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e_{t})-\mathrm{E}[v_{2}(\lambda_{t},e_{t})]$
$=(1-\beta)\big{[}\lambda_{t}u(e_{t}+\tau(e_{t}))-\tau(e_{t})\big{]}+\beta\lambda_{t}\overline{v}_{1}(\lambda_{t})+(1-\beta)\mathrm{E}[\tau(e_{t})],$
and we see that the last terms are maximized by $\tau^{*}$ under the constraint (18). It follows that (16) is written as
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t},t)}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta V(\lambda_{t+1},t+1),$
@@
The task is now to check the terminal condition. Since $u$ is bounded from above, we have
$\lim_{t\to\infty}$ $V(\lambda_{t},t)\beta^{t}=0.$
Thus, $V$ satisfies Terminal condition (i) and from Theorem 1 of Wiszniewska-Matyszkiel [2011], we see $V$ defined in (14) is the value function of the Bellman equation in (15), and the control function $\tau^{*}$ maximises $J$ in (2). ∎
7 Conclusion
This paper presents an alternative solution to avoid the prediction of the previous work: if the incomes of infinitely lived individuals are unobservable, efficient allocations are achieved only at the cost of invoking permanent inequality, leading to an immiserization of society. The proposed inter-period transfer mechanism achieves within-period full insurance by postponing risks. It does not trade off equal opportunities to become wealthy. It could be sustained by intergenerational cooporation. The result sheds light on efficient resource allocation for sustainable societies with equal opportunities.

### v15

@@ Appendix E Proof of Proposition 1
We observe first that if (8) holds, then $V$ defined in (11) satisfies DP. Indeed, since by definition of $\tau^{*}$ we have $u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))=\overline{v}_{1}(\lambda_{t})$,
$V(\lambda_{t},t)$ $=\mathbb{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}\big{[}\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}\big{[}\mathbb{E}_{t+1}\big{[}\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)].$ (E.1)
But $\tau^{*}$ is the solution to the problem
$\sup_{\tau\in\Lambda(\lambda_{t})}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)].$ (E.2)
@@
Indeed, for given information in period $t$, the integrand of the objective function in (E.2) multiplied by $\lambda_{t}$ is written as
$\lambda_{t}\big{[}(1-\beta)$ $u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}$
$=\lambda_{t}\big{[}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e_{t})-\overline{v}_{2}(\lambda_{t})$
and we see that the last terms are maximized by $\tau^{*}$. It follows that (E.1) is written as
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)],$
@@
The task is now to check the terminal condition. Since $u$ is bounded from above, $\lim_{t\to\infty}V(\lambda_{t},t)\beta^{t}=0$. Hence, the sufficient terminal condition of Wiszniewska-Matyszkiel [2011:  Thm.1] is met, and we see $V$ defined in (11) is the value function of the Bellman equation in DP, and $\tau^{*}$ maximizes $J$ in (2). ∎